# Towards LLM Unlearning Resilient to Relearning Attacks: A Sharpness-Aware Minimization Perspective and Beyond

### v3

@@ 5.1 Experiment setups
Datasets and models.
To showcase the robustness improvements brought by SAM and other smoothing techniques, we perform experiments on two representative benchmarks: (1) WMDP (Li et al., 2024), as used in Fig. 1, which evaluates the unlearning capability in hazardous domains, such as biosecurity, cybersecurity, and chemical security. Our experiments primarily focus on the biosecurity aspect of WMDP; (2) MUSE (Shi et al., 2024), which features two distinct unlearning scenarios: forgetting text segments from the Harry Potter book series (labeled ‘Books’) and forgetting news articles from BBC News (labeled ‘News’). Following the literature, we use Zephyr-7B-beta as the original model for WMDP, LLaMA-2 7B fine-tuned on BBC news for News, and ICLM 7B fine-tuned on Harry Potter books for Books. These models, prior to unlearning, are referred to as ‘Origin’, consistent with the terminology in Fig. 1.
LLM unlearning methods and evaluation.
For the WMDP benchmark, we use NPO (Zhang et al., 2024a) with retain regularization as the primary unlearning baseline, as formulated by (2). Additionally, we include representation misdirection for unlearning (RMU) (Li et al., 2024) and gradient difference (GradDiff) (Maini et al., 2024; Liu et al., 2022a) as supplementary baselines. For MUSE, we adopt NPO as the baseline due to its state-of-the-art (SOTA) performance on this benchmark (Shi et al., 2024). More implementation details are provided in Appendix C.
@@
Smoothness optimization implementation. We integrate SAM, RS, CR, GP, and WA with LLM unlearning. For SAM, we set the perturbation parameter $\rho=0.01$ in (6). Table A3 in Appendix E presents a sensitivity study on $\rho$. We find that when $\rho$ is too small (e.g., 0.001), SAM provides limited improvement against relearning attacks. Conversely, when $\rho$ is too large (e.g., 0.1), the perturbations hinder unlearning effectiveness. Additional smoothness optimization details can be found in Appendix C.
@@ 5.2 Experiment results
Evaluation on SAM-integrated unlearning methods beyond NPO.
In Fig. 3, we show the applicability and effectiveness of SAM when integrated with multiple unlearning methods, including NPO, GradDiff (Maini et al., 2024), and RMU (Li et al., 2024). As we can see, all SAM-based variants enhance the robustness of their non-SAM counterparts against relearning attacks. Notably, this improvement does not compromise UT or UE in the absence of relearning attacks. The detailed UE and UT are provided in Table A1 of Appendix D. RMU-type methods achieve better UT (0.57) compared to NPO or GradDiff-type methods (around 0.45). However, they exhibit weaker robustness against relearning attacks compared to NPO+SAM. This discrepancy arises because RMU achieves unlearning by updating only a subset of the model parameters (layers 5, 6, and 7) to balance unlearning with utility preservation. By contrast, relearning attacks can target the entire model, leading to a mismatch in parameter updates that may compromise RMU’s robustness. In Fig. A2 of Appendix D , we further analyze the relationship between the number of parameters involved in smoothness optimization and unlearning robustness by examining RMU.
@@
Unlearning robustness vs. relearning attacks with different relearning epoch counts and data amounts.
In Table. 1, we showcase the UE of NPO and its smoothness optimization-based variants (integrated with SAM, RS, GP, CR, and WA) on WMDP, against the varying number of epochs ($M$) and the forget data amount ($N$) used in relearning attacks. As we can see, UE decreases as either $M$ or $N$ increases. However, compared to the vanilla NPO approach, which nearly reverts to pre-unlearning performance (i.e., ‘Origin’ in Fig. 3) under relearning attacks with $M\geq 2$ and $N\geq 40$, all proposed smooth variants of NPO exhibit much better robustness. Among these, NPO+SAM consistently outperforms the others, demonstrating the strongest resilience against relearning attacks. Additionally, compared to increasing the number of relearning epochs, using a larger number of forget data samples for relearning leads to a more rapid decline in unlearning effectiveness.
@@
Unlearning robustness over diverse relearn sets.
Fig. 4 illustrates the robustness of unlearning against relearning attacks using datasets (AGNews, GSM8K, and SST2) as motivated by (Łucki et al., 2024). As shown, the UE of NPO+SAM after the relearning attacks consistently outperforms that of vanilla NPO. This suggests that, beyond the relearning attacks on the forget set, the robustness of the unlearned model using NPO+SAM generalizes to various types of relearning attacks, even when the relearn sets are derived from datasets different from the forget set.
@@
Evaluation on MUSE dataset.
Fig. 5 compares the unlearning robustness of NPO with NPO + SAM on the MUSE Books and MUSE News datasets. Recall that unlearning effectiveness on MUSE is evaluated using knowledge memorization (KnowMem) and verbatim memorization (VerbMem) on the forget set $\mathcal{D_{\mathrm{f}}}$, with lower values indicating better unlearning effectiveness. As we can see, under relearning attacks with varying numbers of relearn samples (75, 100, 125), NPO+SAM consistently improves the robustness of NPO, as evidenced by lower KnowMem and VerbMem values. Furthermore, changes in VerbMem on $\mathcal{D}_{\mathrm{f}}$ after the relearning attacks are more pronounced compared to those in KnowMem on $\mathcal{D}_{\mathrm{f}}$. This indicates that unlearning precise tokens (VerbMem) is more vulnerable to relearning attacks than unlearning general knowledge encoded in the tokens (KnowMem). In addition to UE, utility performance results are provided in Table A2 in Appendix E.
@@
Unlearning robustness against jailbreaking attacks.
In Fig. 6-(a), we present the unlearning effectiveness of NPO and its smooth enhancements on WMDP under (input-level) adversarial prompts generated by the enhanced GCG (Łucki et al., 2024). As we can see, NPO+SAM and NPO+RS yield lossless UE under jailbreaking attacks, while NPO suffers a significant drop in UE. This is because NPO+SAM and NPO+RS introduce weight smoothing through worst-case and randomized perturbations, respectively. These smoothing effects are known to be helpful in defending against input-level adversarial attacks (Xu et al., 2022; Wei et al., 2023; Zhang et al., 2024b; Cohen et al., 2019). We also provide generation examples under jailbreaking attacks for NPO and NPO+SAM in Table A4 of Appendix F. Thus, our proposal improves resistance to not only relearning attacks (which perturb model weights) but also jailbreaking attacks (which perturb input prompts).
@@
In Fig. 6-(b), we further investigate why smoothness optimization improves robustness against jailbreaking attacks by plotting the KL divergence between the unlearned model and the original model for each output token. A higher KL divergence indicates more effective unlearning. As we can see, the KL divergence for NPO at the first few tokens is notably small, suggesting insufficient unlearning for these ‘shallow’ tokens. This aligns with the known ‘shallow alignment issue’, a limitation in model alignment against jailbreaking attacks (Qi et al., 2025). In contrast, the use of smoothness optimization alleviates this issue, as the first few tokens are effectively unlearned. This improvement explains the enhanced robustness of smoothness optimization against jailbreaking attacks.
@@ A Algorithm for SAM-enhanced Unlearning
1:Original model ${\bm{\theta}}$, forget set $\mathcal{D}_{\mathrm{f}}$, retain set $\mathcal{D}_{\mathrm{r}}$, unlearning steps $N$, learning rate $\eta$, perturbation radius $\rho$, retain regularization $\lambda$
2:${\bm{\theta}}_{\mathrm{u}}\leftarrow{\bm{\theta}}$
3:for $i=1$ to $N$ do
4:     Sample $(x_{\mathrm{f}},y_{\mathrm{f}})\sim\mathcal{D}_{\mathrm{f}}$
5:     $\bm{\delta}\leftarrow\rho\cdot\dfrac{\nabla_{{\bm{\theta}}}\,\ell_{\mathrm{f}}\bigl{(}{\bm{\theta}}_{\mathrm{u}};(x_{\mathrm{f}},y_{\mathrm{f}})\bigr{)}}{\bigl{\|}\nabla_{{\bm{\theta}}}\,\ell_{\mathrm{f}}\bigl{(}{\bm{\theta}}_{\mathrm{u}};(x_{\mathrm{f}},y_{\mathrm{f}})\bigr{)}\bigr{\|}_{2}}$
6:     $g_{\mathrm{f}}\leftarrow\nabla_{{\bm{\theta}}}\,\ell_{\mathrm{f}}\bigl{(}{\bm{\theta}}_{\mathrm{u}}+\bm{\delta};(x_{\mathrm{f}},y_{\mathrm{f}})\bigr{)}$
7:     Sample $(x_{\mathrm{r}},y_{\mathrm{r}})\sim\mathcal{D}_{\mathrm{r}}$
8:     $g_{\mathrm{r}}\leftarrow\nabla_{{\bm{\theta}}}\,\ell_{\mathrm{r}}\bigl{(}{\bm{\theta}}_{\mathrm{u}};(x_{\mathrm{r}},y_{\mathrm{r}})\bigr{)}$
9:     ${\bm{\theta}}_{\mathrm{u}}\leftarrow{\bm{\theta}}_{\mathrm{u}}-\eta\bigl{(}g_{\mathrm{f}}+\lambda\cdot g_{\mathrm{r}}\bigr{)}$
10:end for
11:return ${\bm{\theta}}_{\mathrm{u}}$
@@ C Detailed Experiment Setups
For WMDP (Li et al., 2024), we utilize Zephyr-7B-beta as the original model specified in the benchmark. The dataset includes a forget set composed of plain texts related to biosecurity knowledge and a retain set of unrelated general content from Wikitext (Merity et al., 2016). We perform 125 unlearning steps for both NPO and GradDiff, using grid searches over the learning rate in [$2.5\times 10^{-6}$, $10^{-5}$] and $\lambda$ in [1, 2.5]. For NPO, we additionally tune $\beta$ in [0.01, 0.05]. For RMU, following (Li et al., 2024), we conduct 150 unlearning steps with a grid search for $\lambda$ in the range [800, 1600]. Regarding smoothing methods, we run grid searches for $\rho$ within the range [$10^{-3}$, $10^{-1}$] under NPO + SAM/RS, and $\gamma$ in the range [1, 10] under NPO + CR/GP. In NPO + SWA, we apply model averaging starting at 100 steps and repeating every five steps thereafter. We set the number of perturbation samples for RS to 3. For RMU+SAM, we unlearn in layers 5 to 7 and apply perturbations to layers 1 to 7.
@@ D Additional Results on WMDP
The relationship between robustness and parameter count in smoothness optimization. In Fig. A2, we illustrate the impact of parameter count in smoothness optimization on the loss landscape over $\mathcal{D}_{\mathrm{f}}$ and the unlearning robustness against relearning attacks. RMU is a partial model unlearning method that updates only a subset of layers in the original model, which consists of a total of 32 layers. Fig. A2-(a) presents the vanilla RMU, which performs unlearning at layers 5$\sim$7. It can be observed that its loss landscape undergoes a sharp change at the origin. In contrast, Fig. A2-(b) depicts the SAM-enhanced RMU, which unlearns at layers 5$\sim$7 and applies perturbations at layers 5$\sim$7. As a result, its loss landscape appears slightly smoother compared to Fig. A2-(a). In Fig. A2-(c), the SAM-enhanced RMU not only unlearns at layers 5$\sim$7 but also applies perturbations across layers 1$\sim$7. This results in a smoother loss landscape. Additionally, In Fig. A2-(d) illustrates the unlearning robustness against a relearning attack using 20 samples from the WMDP Bio forget set, trained for 3 epochs. It is evident that as the number of perturbed parameters increases, the model demonstrates greater robustness.

### v4

@@ 5.1 Experiment setups
Datasets and models. To showcase the robustness improvements brought by SAM and other smoothing techniques, we perform experiments on two representative benchmarks: (1) WMDP (Li et al., 2024), as used in Fig. 1, which evaluates the unlearning capability in hazardous domains, such as biosecurity, cybersecurity, and chemical security. Our experiments primarily focus on the biosecurity aspect of WMDP; (2) MUSE (Shi et al., 2024), which features two distinct unlearning scenarios: forgetting text segments from the Harry Potter book series (labeled ‘Books’) and forgetting news articles from BBC News (labeled ‘News’). Following the literature, we use Zephyr-7B-beta and LLaMA-3 8B as the original model for WMDP, LLaMA-2 7B fine-tuned on BBC news for News, and ICLM 7B fine-tuned on Harry Potter books for Books. These models, prior to unlearning, are referred to as ‘Origin’, consistent with the terminology in Fig. 1.
LLM unlearning methods and evaluation. For the WMDP benchmark, we use NPO (Zhang et al., 2024a) with retain regularization as the primary unlearning baseline, as formulated by (2). Additionally, we include representation misdirection for unlearning (RMU) (Li et al., 2024), gradient difference (GradDiff) (Maini et al., 2024; Liu et al., 2022a), RMU with latent adversarial training (RMU-LAT) (Sheshadri et al., 2024) and tampering attack resistance (TAR) (Tamirisa et al., 2024) as supplementary baselines. For MUSE, we adopt NPO as the baseline due to its state-of-the-art performance on this benchmark (Shi et al., 2024). More implementation details are provided in Appendix C.
@@
Smoothness optimization implementation. We integrate SAM, RS, CR, GP, and WA with LLM unlearning. For SAM, we set the perturbation parameter $\rho=0.01$ in (6). Additional smoothness optimization details can be found in Appendix C.
@@ 5.2 Experiment results
Evaluation on SAM-integrated unlearning methods beyond NPO. In Fig. 3, we show the applicability and effectiveness of SAM when integrated with multiple unlearning methods, including NPO, GradDiff (Maini et al., 2024), and RMU (Li et al., 2024). As we can see, all SAM-based variants enhance the robustness of their non-SAM counterparts against relearning attacks. Notably, this improvement does not compromise UT or UE in the absence of relearning attacks. The detailed UE and UT are provided in Table A1 of Appendix D. RMU-type methods achieve better UT (0.57) compared to NPO or GradDiff-type methods ($\sim$0.45). However, they exhibit weaker robustness against relearning attacks compared to NPO+SAM. This discrepancy arises because RMU achieves unlearning by updating only a subset of the model parameters (layers 5, 6, and 7) to balance unlearning with utility preservation. By contrast, relearning attacks can target the entire model, leading to a mismatch in parameter updates that may compromise RMU’s robustness. In Fig. A2 of Appendix D , we further analyze the relationship between the number of parameters involved in smoothness optimization and unlearning robustness by examining RMU.
@@
Methods UT (↑) UE (↑) Time (min) (↓)
W/o atk W/ atk
NPO 0.50 0.73 0.41 5.8
TAR 0.54 0.74 0.70 7441.9
RMU-LAT 0.56 0.70 0.44 10.3
NPO+SAM 0.51 0.74 0.70 11.5
In Table 1, we provide additional comparisons of NPO+SAM against other robust unlearning methods, including TAR (Tamirisa et al., 2024) and RMU-LAT (Sheshadri et al., 2024), evaluated on a different model, LLaMA-3 8B. The results show that NPO+SAM achieves highly competitive performance on the WMDP benchmark, matching TAR and significantly outperforming both the vanilla NPO and RMU-LAT. The strong performance gap between NPO+SAM and RMU-LAT underscores the effectiveness of weight-space perturbations (employed by SAM) over activation-space perturbations (used by RMU-LAT) in defending against relearning. Although TAR approaches the unlearning-versus-relearning problem via a meta-learning framework, its reliance on meta-gradients and multi-step gradient unrolling introduces substantial computational overhead. In contrast, NPO+SAM achieves a superior balance between unlearning efficacy, robustness, and efficiency, offering a more practical and scalable solution.
Table 2: Unlearning robustness comparison of NPO and its smoothness optimization-based variants on WMDP under different relearning attacks settings. $N$ represents the number of forget samples used for relearning with 1 epoch, and $M$ denotes the number of relearning epochs using 20 forget samples. The best robustness in each relearning setting is highlighted in red . The table format is consistent with Table 1.
@@
Unlearning robustness vs. relearning attacks with different relearning epoch counts and data amounts. In Table. 2, we showcase the UE of NPO and its smoothness optimization-based variants (integrated with SAM, RS, GP, CR, and WA) on WMDP, against the varying number of epochs ($M$) and the forget data amount ($N$) used in relearning attacks. As we can see, UE decreases as either $M$ or $N$ increases. However, compared to the vanilla NPO approach, which nearly reverts to pre-unlearning performance (i.e., ‘Origin’ in Fig. 3) under relearning attacks with $M\geq 2$ and $N\geq 40$, all proposed smooth variants of NPO exhibit much better robustness. Among these, NPO+SAM consistently outperforms the others, demonstrating the strongest resilience against relearning attacks. Additionally, compared to increasing the number of relearning epochs, using a larger number of forget data samples for relearning leads to a more rapid decline in unlearning effectiveness.
@@
Unlearning robustness over diverse relearn sets. Fig. 4 illustrates the robustness of unlearning against relearning attacks using datasets (AGNews, GSM8K, and SST2) as motivated by (Łucki et al., 2024). As shown, the UE of NPO+SAM after the relearning attacks consistently outperforms that of vanilla NPO. This suggests that, beyond the relearning attacks on the forget set, the robustness of the unlearned model using NPO+SAM generalizes to various types of relearning attacks, even when the relearn sets are derived from datasets different from the forget set.
@@
Evaluation on MUSE dataset. Fig. 5 compares the unlearning robustness of NPO with NPO + SAM on the MUSE Books and MUSE News datasets. Recall that unlearning effectiveness on MUSE is evaluated using knowledge memorization (KnowMem) and verbatim memorization (VerbMem) on the forget set $\mathcal{D_{\mathrm{f}}}$, with lower values indicating better unlearning effectiveness. As we can see, under relearning attacks with varying numbers of relearn samples (75, 100, 125), NPO+SAM consistently improves the robustness of NPO, as evidenced by lower KnowMem and VerbMem values. Furthermore, changes in VerbMem on $\mathcal{D}_{\mathrm{f}}$ after the relearning attacks are more pronounced compared to those in KnowMem on $\mathcal{D}_{\mathrm{f}}$. This indicates that unlearning precise tokens (VerbMem) is more vulnerable to relearning attacks than unlearning general knowledge encoded in the tokens (KnowMem). In addition to UE, utility performance results are provided in Table A2 in Appendix E.
Unlearning robustness against jailbreaking attacks and its connection to ‘shallow unlearning alignment’ issue. In Fig. 6-(a), we present the unlearning effectiveness of NPO and its smooth enhancements on WMDP under (input-level) adversarial prompts generated by the enhanced GCG (Łucki et al., 2024). As we can see, NPO+SAM and NPO+RS yield lossless UE under jailbreaking attacks, while NPO suffers a significant drop in UE. This is because NPO+SAM and NPO+RS introduce weight smoothing through worst-case and randomized perturbations, respectively. These smoothing effects are known to be helpful in defending against input-level adversarial attacks (Xu et al., 2022; Wei et al., 2023; Zhang et al., 2024b; Cohen et al., 2019). We also provide generation examples under jailbreaking attacks for NPO and NPO+SAM in Table A4 of Appendix F. Thus, our proposal improves resistance to not only relearning attacks (which perturb model weights) but also jailbreaking attacks (which perturb input prompts).
In Fig. 6-(b), we further investigate why smoothness optimization improves robustness against jailbreaking attacks by plotting the KL divergence between the unlearned model and the original model for each output token. A higher KL divergence indicates more effective unlearning. As we can see, the KL divergence for NPO at the first few tokens is notably small, suggesting insufficient unlearning for these ‘shallow’ tokens. This phenomenon aligns with the well-known shallow safety alignment issue, which highlights the limitations of current safety alignment techniques against jailbreaking attacks (Qi et al., 2025). In our context, we refer to this limitation as shallow unlearning alignment. In contrast, the use of smoothness optimization alleviates this issue, as the first few tokens are effectively unlearned. This improvement explains the enhanced robustness of smoothness optimization against jailbreaking attacks.
@@
Ablation study on SAM’s hyperparameter $\rho$. Table A3 in Appendix E presents a sensitivity study on $\rho$. We find that when $\rho$ is too small (e.g., 0.001), SAM provides limited improvement against relearning attacks. Conversely, when $\rho$ is too large (e.g., 0.1), the perturbations hinder unlearning effectiveness.
@@ 6 Conclusion
Impact Statement
Our research enhances the robustness of LLM unlearning against relearning and jailbreaking attacks by leveraging smoothness optimization, thereby strengthening data privacy and regulatory compliance. By integrating techniques such as sharpness-aware minimization (SAM), we achieve more reliable unlearning, reducing unintended knowledge retention and reinforcing model security. Furthermore, this study establishes a critical link between smoothness optimization and unlearning, helping bridge the gap between foundational optimization research and use-inspired advancements in LLM unlearning. However, enhanced unlearning could be misused to selectively erase essential knowledge, while stronger resistance to relearning may hinder the recovery of valuable information. To address these risks, strict ethical standards and regulatory oversight are essential. Future research should prioritize governance, fairness, and auditing to ensure AI technologies are developed responsibly and transparently.
@@ A Algorithm for SAM-enhanced Unlearning
0:  Original model ${\bm{\theta}}$, forget set $\mathcal{D}_{\mathrm{f}}$, retain set $\mathcal{D}_{\mathrm{r}}$, unlearning steps $N$, learning rate $\eta$, perturbation radius $\rho$, retain regularization $\lambda$.
1:  ${\bm{\theta}}_{\mathrm{u}}\leftarrow{\bm{\theta}}$
2:  for $i=1$ to $N$ do
3:     Sample $(x_{\mathrm{f}},y_{\mathrm{f}})\sim\mathcal{D}_{\mathrm{f}}$
4:     $\bm{\delta}\leftarrow\rho\cdot\dfrac{\nabla_{{\bm{\theta}}}\,\ell_{\mathrm{f}}\bigl{(}{\bm{\theta}}_{\mathrm{u}};(x_{\mathrm{f}},y_{\mathrm{f}})\bigr{)}}{\bigl{\|}\nabla_{{\bm{\theta}}}\,\ell_{\mathrm{f}}\bigl{(}{\bm{\theta}}_{\mathrm{u}};(x_{\mathrm{f}},y_{\mathrm{f}})\bigr{)}\bigr{\|}_{2}}$
5:     $g_{\mathrm{f}}\leftarrow\nabla_{{\bm{\theta}}}\,\ell_{\mathrm{f}}\bigl{(}{\bm{\theta}}_{\mathrm{u}}+\bm{\delta};(x_{\mathrm{f}},y_{\mathrm{f}})\bigr{)}$
6:     Sample $(x_{\mathrm{r}},y_{\mathrm{r}})\sim\mathcal{D}_{\mathrm{r}}$
7:     $g_{\mathrm{r}}\leftarrow\nabla_{{\bm{\theta}}}\,\ell_{\mathrm{r}}\bigl{(}{\bm{\theta}}_{\mathrm{u}};(x_{\mathrm{r}},y_{\mathrm{r}})\bigr{)}$
8:     ${\bm{\theta}}_{\mathrm{u}}\leftarrow{\bm{\theta}}_{\mathrm{u}}-\eta\bigl{(}g_{\mathrm{f}}+\lambda\cdot g_{\mathrm{r}}\bigr{)}$
9:  end for
10:  return ${\bm{\theta}}_{\mathrm{u}}$
@@ C Detailed Experiment Setups
For WMDP (Li et al., 2024), we utilize Zephyr-7B-beta as the original model specified in the benchmark. The dataset includes a forget set composed of plain texts related to biosecurity knowledge and a retain set of unrelated general content from Wikitext (Merity et al., 2016). We perform 125 unlearning steps for both NPO and GradDiff, using grid searches over the learning rate in [$2.5\times 10^{-6}$, $10^{-5}$] and $\lambda$ in [1, 2.5]. For NPO, we additionally tune $\beta$ in [0.01, 0.05]. For RMU, following Li et al. (2024), we conduct 150 unlearning steps with a grid search for $\lambda$ in the range [800, 1600]. Regarding smoothing methods, we run grid searches for $\rho$ within the range [$10^{-3}$, $10^{-1}$] under NPO + SAM/RS, and $\gamma$ in the range [1, 10] under NPO + CR/GP. In NPO + SWA, we apply model averaging starting at 100 steps and repeating every five steps thereafter. We set the number of perturbation samples for RS to 3. For RMU+SAM, we unlearn in layers 5 to 7 and apply perturbations to layers 1 to 7.
@@ D Additional Results on WMDP
The relationship between robustness and parameter count in smoothness optimization. In Fig. A2, we illustrate the impact of parameter count in smoothness optimization on the loss landscape over $\mathcal{D}_{\mathrm{f}}$ and the unlearning robustness against relearning attacks. Fig. A2-(a) presents the vanilla RMU, which performs unlearning at layers 5$\sim$7. It can be observed that its loss landscape undergoes a sharp change at the origin. In contrast, Fig. A2-(b) depicts the SAM-enhanced RMU, which unlearns at layers 5$\sim$7 and applies perturbations at layers 5$\sim$7, with the perturbation weights accounting for 2.43% of the total model parameters. As a result, its loss landscape appears slightly smoother compared to Fig. A2-(a). In Fig. A2-(c), the SAM-enhanced RMU not only unlearns at layers 5$\sim$7 but also applies perturbations across layers 1$\sim$7, with perturbation weights making up 5.68% of the total model parameters. This results in a smoother loss landscape. Additionally, In Fig. A2-(d) illustrates the unlearning robustness against a relearning attack using 20 samples from the WMDP Bio forget set, trained for 3 epochs. It is evident that as the number of perturbed parameters increases, the model demonstrates greater robustness.